# Dynamics of Moral Behavior in Heterogeneous Populations of Learning Agents

### v7

@@ Introduction
Our environment involves a population $P$ (of size $N=16$) playing an iterated game. At each iteration, a moral player $M$ and an opponent $O$ play a simultaneous one-shot Prisoner’s Dilemma game with two possible actions - Cooperate or Defect ($a^{t}_{M},a^{t}_{O}\in\{C,D\}$, see Table 1 for the respective payoffs). To allow learning in a population, we run our simulation in 30000 episodes, where a single episode involves each player selecting a partner once, and then each corresponding pair of players playing the game. For clarity, each agent gets to make a selection on each episode, so $N$ games are played. In particular, at iteration $t$, the learning agent $M$ observes a selection state including the action played by each possible opponent $O\in P$ at $t-1$: $s_{M,sel}^{t}=[a_{O_{1}}^{t-1},...,a_{O_{n-1}}^{t-1}]$. Using this state and the current learned policy, $M$ selects an opponent $O$. Then, to choose an action for the dilemma, player $M$ relies on the state representing the latest move of their selected opponent $O$: $s_{M,dil}^{t}=a_{O}^{t-1}$. Simultaneously, their opponent $O$ chooses an action following the same principle. The players $M$ and $O$ then each receive a game reward $R_{M}^{t+1}$ and $R_{O}^{t+1}$ (corresponding to the game payoffs) and observe a new state $s_{M,dil}^{t+1}$ and $s_{O,dil}^{t+1}$ based on their opponent’s move $a^{t}$.
@@
Pro-Social De Deontological $R_{M_{intr}}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise}\end{cases}$ External norm
@@
V-Ki Virtue-Kindness $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=C\\ 0,&\text{otherwise}\end{cases}$ Internal norm
@@
Anti-Social mDe Malicious Deontological $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise}\end{cases}$ External norm
@@
V-Ag Virtue-Aggression $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=D\\ 0,&\text{otherwise}\end{cases}$ Internal norm
@@
$\pi(s_{t})=\begin{cases}\arg\max_{a}Q(s_{t},a),&\text{with probability }1-\epsilon\\ U(A=\{C,D\}),&\text{with probability }\epsilon.\end{cases}$ (2)
@@
  • the Deontological (De) agent tries to follow the norm of conditional cooperation (Kant 1785; Fehr and Fischbacher 2004) and gets penalized through the negative reward $-\xi$ for defecting against a cooperator (i.e., an opponent who previously cooperated);
@@ Results
The behavior of the majority-De population is the most striking. Here, aUt and V-Ag players are relatively popular. A further analysis of the specific selection patterns among individual players in this population (Figure 4b) reveals that the De players in particular tend to mostly prefer anti-social opponents aUt and V-Ag. This apparently self-sabotaging behavior can be explained by a closer analysis of the De player’s $R_{intr}$ they are penalized for defecting against a cooperator, so the safest behavior in a heterogeneous population is to always select a defector, so that they have the lowest chance of violating their moral norm. Figure 4b also shows that another player - aUt - prefers to avoid the De majority players and instead tends to select the V-Ag opponent. This highlights that minimizing collective reward (which is the preference embedded in the aUt agent’s intrinsic reward) is best achieved by avoiding the very cooperative De agent. Thus, we find an interesting dynamic in the heterogeneous majority-De population, in which the majority of the players follow a strong cooperative norm, demonstrating ways in which these De players can be very cooperative (see Figure 1a and Appendix, Figure 6) but not very popular.
@@ Acknowledgments
Elizaveta Tennant was supported by the Leverhulme Trust (DS-2017-026; Doctoral Training Programme for the Ecological Study of the Brain).
@@ Appendix D Partner Selection Algorithm
The algorithm for partner selection ($sel$) and at the basis of the dilemma game ($dil$) with 2 separate Q-network models is presented below (see Algorithm 1). We use episodes, so that all players learn simultaneously, after each player had a single chance to select and play. For clarity, each selector’s state $s_{i,sel}^{t}$ relies on the same $environment\_state$ (i.e., simultaneous selections), and therefore every player $i$’s $s_{i,sel}^{t+1}$ in the $sel$ trajectory only differs from $s_{i,sel}^{t}$ by one entry: $a_{j,dil}^{t}$ at player $j$’s position. For $dil$, each player $i$ uses all of their observed dilemma trajectories to update their network, by calculating an average loss across these trajectories.
@@
0:  population of players $P$ of size $N$, total episodes $T=30000$
1:  set random initial $environment\_state$ (i.e., one random move for each player)
@@
9:     Get current global $environment\_state$ from the environment
@@
15:           $i$ observes the $environment\_state$ and removes itself to parse in their own $s_{i,sel}^{t}$
@@
26:           $i$ and $j$ parse states $s_{i,dil}^{t}$, $s_{j,dil}^{t}$ from the $environment\_state$ the current opponent’s index
@@
30:        $i$ and $j$ receive rewards $r_{i,dil}^{t}$ $r_{j,dil}^{t}$ respectively
@@
36:     PART 3: Update $sel$ and $dil$ networks
@@
38:        update the corresponding part of $environment\_state$ with the latest moves by $i$ ($a_{i,dil}^{t}$)
39:        player $i$ observes their $s_{i,sel}^{t+1}$ by copying the latest $environment\_state$ and removing self
@@
41:        player $i$ samples one experience from $D_{i,sel}$: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$ calculates loss
42:        player $i$ samples all experiences from $D_{i,dil}$: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$ calculates average loss
43:        player $i$ updates $Q_{i,sel}$ network with respect to time step $t$ the calculated loss
44:        player $i$ updates $Q_{i,dil}$ network with respect to time step $t$ the calculated loss

### v8

@@ Introduction
Our environment involves a population $P$ (of size $N=16$) playing an iterated game. At each iteration, a moral player $M$ and an opponent $O$ play a simultaneous one-shot Prisoner’s Dilemma game with two possible actions - Cooperate or Defect ($a^{t}_{M},a^{t}_{O}\in\{C,D\}$, see Table 1 for the respective payoffs). To allow learning in a population, we run our simulation in 30000 episodes, where a single episode involves each player selecting a partner once, and then each corresponding pair of players playing the game. For clarity, each agent gets to make a selection on each episode, so $N$ games are played. In particular, at iteration $t$, the learning agent $M$ observes a selection state including the action played by each possible opponent $O\in P$ at t1t-1: $s_{M,sel}^{t}=[a_{O_{1}}^{t-1},...,a_{O_{n-1}}^{t-1}]$. Using this state and the current learned policy, $M$ selects an opponent $O$. Then, to choose an action for the dilemma, player $M$ relies on the state representing the latest move of their selected opponent $O$: $s_{M,dil}^{t}=a_{O}^{t-1}$. Simultaneously, their opponent $O$ chooses an action following the same principle. The players $M$ and $O$ then each receive a game reward $R_{M}^{t+1}$ and $R_{O}^{t+1}$ (corresponding to the game payoffs) and observe a new state $s_{M,dil}^{t+1}$ and $s_{O,dil}^{t+1}$ based on their opponent’s move $a^{t}$.
@@
Pro-Social De Deontological $R_{M_{intr}}^{t}=\begin{cases}$--$\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise}\end{cases}\$External norm
@@
V-Ki Virtue-Kindness $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=C\\ 0,&\text{otherwise}\end{cases}\$Internal norm
@@
Anti-Social mDe Malicious Deontological $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=D,a_{O}^{t-1}=C\\ 0,&\text{otherwise}\end{cases}\$External norm
@@
V-Ag Virtue-Aggression $R_{M_{intr}}^{t}=\begin{cases}\xi,&\text{if }a_{M}^{t}=D\\ 0,&\text{otherwise}\end{cases}\$Internal norm
@@
$\pi(s_{t})=\begin{cases}\arg\max_{a}Q(s_{t},a),&\text{with probability }1-\epsilon\\ U(A=\{C,D\}),&\text{with probability }\epsilon.\end{cases}\$(2)
@@
  • the Deontological (De) agent tries to follow the norm of conditional cooperation (Kant 1785; Fehr and Fischbacher 2004) and gets penalized through the negative reward $\xi$ for defecting against a cooperator (i.e., an opponent who previously cooperated);
@@ Results
The behavior of the majority-De population is the most striking. Here, aUt and V-Ag players are relatively popular. A further analysis of the specific selection patterns among individual players in this population (Figure 4b) reveals that the De players in particular tend to mostly prefer anti-social opponents aUt and V-Ag. This apparently self-sabotaging behavior can be explained by a closer analysis of the De player’s $R_{intr}$ - they are penalized for defecting against a cooperator, so the safest behavior in a heterogeneous population is to always select a defector, so that they have the lowest chance of violating their moral norm. Figure 4b also shows that another player - aUt - prefers to avoid the De majority players and instead tends to select the V-Ag opponent. This highlights that minimizing collective reward (which is the preference embedded in the aUt agent’s intrinsic reward) is best achieved by avoiding the very cooperative De agent. Thus, we find an interesting dynamic in the heterogeneous majority-De population, in which the majority of the players follow a strong cooperative norm, demonstrating ways in which these De players can be very cooperative (see Figure 1a and Appendix, Figure 6) but not very popular.
@@ Acknowledgments
Elizaveta Tennant was supported by the Leverhulme Trust (DS-2017-026; Doctoral Training Programme for the Ecological Study of the Brain). Mirco Musolesi was supported by the Italian Ministry of University and Research (MUR) through the project PRIN 2022 “Machine-learning based control of complex multi-agent systems for search and rescue operations in natural disasters (MENTOR)” funded by the European Union - NextGenerationEU.
@@ Appendix D Partner Selection Algorithm
The algorithm for partner selection (selsel) and at the basis of the dilemma game (dildil) with 2 separate Q-network models is presented below (see Algorithm 1). We use episodes, so that all players learn simultaneously, after each player had a single chance to select and play. For clarity, each selector’s state $s_{i,sel}^{t}$ relies on the same environment_stateenvironment\_state (i.e., simultaneous selections), and therefore every player $i$’s $s_{i,sel}^{t+1}$ in the selsel trajectory only differs from $s_{i,sel}^{t}$ by one entry: $a_{j,dil}^{t}$ at player $j$’s position. For dildil, each player $i$ uses all of their observed dilemma trajectories to update their network, by calculating an average loss across these trajectories.
@@
0: population of players $P$ of size NN, total episodes $T=30000$
1: set random initial environment_stateenvironment\_state (i.e., one random move for each player)
@@
9:  Get current global environment_stateenvironment\_state from the environment
@@
15:    $i$ observes the environment_stateenvironment\_state and removes itself to parse in their own $s_{i,sel}^{t}$
@@
26:    $i$ and $j$ parse states $s_{i,dil}^{t}$, $s_{j,dil}^{t}$ from the environment_stateenvironment\_state & the current opponent’s index
@@
30:   $i$ and $j$ receive rewards $r_{i,dil}^{t}$ & $r_{j,dil}^{t}$ respectively
@@
36:  PART 3: Update selsel and dildil networks
@@
38:   update the corresponding part of environment_stateenvironment\_state with the latest moves by $i$ ($a_{i,dil}^{t}$)
39:   player $i$ observes their $s_{i,sel}^{t+1}$ by copying the latest environment_stateenvironment\_state and removing self
@@
41:   player $i$ samples one experience from $D_{i,sel}$: $(s_{i,sel}^{t},a_{i,sel}^{t},r_{i,sel}^{t},s_{i,sel}^{t+1})$ & calculates loss
42:   player $i$ samples all experiences from $D_{i,dil}$: $(s_{i,dil}^{t},a_{i,dil}^{t},r_{i,dil}^{t},s_{i,dil}^{t+1})$ & calculates average loss
43:   player $i$ updates $Q_{i,sel}$ network with respect to time step $t$ & the calculated loss
44:   player $i$ updates $Q_{i,dil}$ network with respect to time step $t$ & the calculated loss